### v5

@@ IV Future opportunities
In this review, we examined publications that employed SLAM techniques in their navigation approaches. One of the distinct advantages of SLAM is its applicability in diverse locations without the need for pre-built maps or additional infrastructure such as Bluetooth beacons or RFID tags. However, there is still room for improvement in various aspects of these systems, including their ability to handle complex scenarios, provide accurate obstacle information, and seamlessly transition between indoor and outdoor environments. This section discusses the open problems and research directions identified during the SLR.

### v6

@@ IV Future opportunities
In this review, we examined publications that employed SLAM techniques in their navigation approaches. One of the distinct advantages of SLAM is its applicability in diverse locations without the need for pre-built maps or additional infrastructure such as Bluetooth beacons or RFID tags. However, there is still room for improvement in various aspects of these systems, including their ability to handle complex scenarios, provide accurate obstacle information, and seamlessly transition between indoor and outdoor environments.
There is a notable lack of focus on adapting SLAM to challenging situations especially dynamic environments, specifically for visually impaired navigation. Studies in this field typically employ SLAM as a pre-existing tool or use it without any significant adaptation to effectively handle challenges. Researchers in this area can draw inspiration from recent advancements in robotics to address this gap. By leveraging cutting-edge techniques from robotic navigation, SLAM systems that are more robust and suitable for assisting visually impaired individuals in real-world scenarios can be developed.
This section discusses the open problems and research directions identified during the SLR.
@@
Human-computer interaction for SLAM-based assistive devices An important area for future research is to improve the human-computer interaction (HCI) aspects of SLAM-based assistive devices for BVI individuals. Although SLAM techniques have shown great potential in gathering and processing environmental information, effectively communicating this information to BVI users remains a significant challenge. Future research should focus on developing intuitive and non-intrusive methods to convey complex spatial data and navigational instructions to BVI users. This includes:
  • Multi-modal feedback systems: Exploring combinations of audio, haptic, and other non-visual feedback methods to provide rich contextual information without overwhelming the user.
  • Adaptive interfaces: Developing interfaces that can adjust the level and type of information provided based on the user’s preferences, familiarity with the environment, and the current situation.
  • Natural language processing: Improving the ability of systems to understand and respond to natural language queries, allowing for a more intuitive interaction between the user and device.
  • Cognitive load optimization: Investigating ways to balance the provision of detailed environmental information, ensuring that users receive the necessary guidance without cognitive overload.
  • Real-time situational awareness: Developing methods to effectively communicate dynamic elements of the environment, such as moving obstacles or changing traffic conditions in real-time.
Addressing these HCI challenges will be crucial in translating the technical capabilities of SLAM into practical, user-friendly assistive devices that can significantly enhance the mobility and independence of BVI individuals. Future research in this area should involve close collaboration with BVI users to ensure that the developed interfaces meet their needs and preferences.